In a 300-m race, runner A starts from rests and accelerates at 1.6m/s2


for the first 30 m and then runs


at a constant speed. Runner B starts from rests and accelerates at 2.0m/s2


for the first 30m and then

runs at a constant speed. Runner A begins running as soon as the race begins but B firsts takes a nap

to rest up. What is the longest nap that B can take and still not to lose the race?


1
Expert's answer
2022-03-14T13:13:29-0400

Solution.

"S=300m;"

"v_{0A}=0;"

"a_A=1.6 m\/s^2;"

"v_{0B}=0;"

"a_B=2.0 m\/s^2;"

"S_A=S_B=30m;"

"S_a=\dfrac{v_A^2}{2a_A}\implies v_A=\sqrt{2a_AS_A};"

"v_A=\sqrt{2\sdot1.6\sdot30}=9.8 m/s;"

"a_A=\dfrac{{v_A-v_{0A}}}{t_A}\implies t_A=\dfrac{v_A-v_{0A}}{a_A};"

"t_A=\dfrac{9.8}{1.6}=6.1s;"

"S_B=\dfrac{v_B^2}{2a_B}\implies v_B=\sqrt{2a_BS_B};"

"v_B=\sqrt{2\sdot2.0\sdot30}=11.0 m/s;"

"a_B=\dfrac{{v_B-v_{0B}}}{t_B}\implies t_B=\dfrac{v_B-v_{0B}}{a_B};"

"t_B=\dfrac{11.0}{2.0}=5.5s;"


"\tau_A=\dfrac{270}{9.8}=27.6 s;"


"\tau_B=\dfrac{270}{11.0}=24.5 s;"

"\Delta t=(t_A+\tau_A)-(t_B+\tau_B);"

"\Delta t=(6.1+27.6)-(5.5+24.5)=3.7s;"

Answer: "\Delta t=3.7s."